Questions 2: A) Last summer, Molly had so much fun studying abroad and eating pretzels in Germany that she now wants to buy her own German pretzel bakery. Molly found a pretzel bakery for sale, but Jessica, the current owner, cannot sell it for 7 years. Jessica will accept $68,000 for the bakery in 7 years. Molly just invested $5,000 today at 6%, which she will use to partially pay Jessica. Molly needs to know how much more to invest, 4 years from now at 8%, in order to have enough money to purchase the bakery at the end of 7 years.

B) Assume that after you graduate from college, you are able save $2,000 per year. If you put the money in a 4% interest bearing savings account at the end of each year for 15 years, how much money would you have?

C) Assume you will retire in 20 years and live 20 more years after that. Your annual spending requirements during retirement will be $90,000 per year at the end of each year. If the interest rate is 10%, how much must you invest today to fund your retirement?
Answer #1

A)  

Initial investment = 5000

Interest @ 6% p.a.

Amount to be received after 7 year from this investment =

Amount = P × (1 + R)⁷

= 5000 × (1 + 0.06)⁷

= 7518.15

* Additional amount required to purchase backery

= 68000 - 7518.15

= 60418.85

Additional amount will be invested at the end of 4th year =  

Amount = P ×(1+R)³

60418.85 = P (1.08)³

P = 60418.85 ÷ 1.259712

P = 48012.56

Amount to be invested after 4 years is 48012.56